a rectangle has a perimeter of 160' and the width is x and the lenght is 2x-10. What would the area also be.
Answer by checkley77(12844)

2W+2L=perimeter.
2x+2(2x-10)=160
2x+4x-20=160
6x=160+20
6x=180
x=180/6
x=30
2*30=60 for the width.
2*30-20=60-20=40 for the length.
Area=LW
Area=60*40=2,400 ans.